Class Managernieht Help April 7 Begin Date: 4/5/2017 12:00:00 AM Due Date: 4/7/2017 2:00:00 AM End Date: 4/7/2017 2:00:00 AM (20%) Problem 3: Consider a sheet of paper 8.45 in by 11.5 in. Go How much force, in newtons, is exerted on one side of the paper by the atmosphere sino coso tano 7 8 9 HoME 4 5 6 cotan 1 2 3 atano acotano si cosh00 tanho cotanh0 IVO BACKSPACE CLEAR Degrees O Radians Submit Hint Feedback give up! Hints 1% deduction per hint. Hints remaining: 1 Feedback: 1% deduction per feedback. Grade Summary Deductions 5% 95% Potential Submissions Attempts remaining: 4 50% per attempt) detailed view 5%
Answer #1


length of the paper l = 11.5 in = 11.5*0.0254 = 0.292 m

width of the paper b = 8.45 in = 8.45*0.0254 = 0.215 m


area of the paper A = l*b = 0.292*0.215 = 0.0628 m^2

atmospheric pressure P = 1.013*10^5 Pa

force F = P*A

F = 1.013*10^5*0.0628


F = 6361.64 N<<<----------answer